I need an equation on this problem. Thanks!
What is the sum of 1+2+3+...+20?

Answer by stanbon(75887)

What is the sum of 1+2+3+...+20?
----------
It is an arithmetic sequence.
The sum of n terms is S(n)= (n/2)(1st + last)
Your Problem:
S(20) = (20/2)(1+20)
= 10 *21
= 210
=============
Cheers,
Stan H.